Johnson & Johnson (JNJ) has a forward P/E of 16.26, a Price/Sales of 3.80, a Price/Book of 3.84, a PEG ratio of 2.75, and an enterprise value/EBITDA of 11.40. What is the forecasted growth rate of earnings over the next five years?

A.

3.8%

B.

4.2%

C.

5.9%

D.

4.9%
Answer #1

PEG Ratio = (P/E) / EPS Growth Rate

2.75 = 16.26 / EPS Growth Rate

EPS Growth Rate = 16.26 / 2.75 = 5.9%

Hence, Option "C" is correct.